prove that in a triangle ABC summation(sinAsin(B-C))=0

Grade:

2 Answers

Karthik Pasupulatei
8 years ago
first write sin2A then 2SINA.COSA then write 2SINA.COS(90-B-C) THEN U WRITE 2SINA SIN(B-C)
Lab Bhattacharjee
8 years ago
B+C=\pi-A\implies\sin(B+C)=\cdots=\sin A \\ \sin A\cdot\sin(B-C)=\sin(B+C)\sin(B-C)=\sin^2B-\sin^2C